OpenStudy (osanseviero):

How can i solve this? 2(x-2)(x+1)>0

OpenStudy (osanseviero):

The original question was: x(3x+2)>(x+2)^2 so 3(x^2)+2x>(x^2)+4x+4 so: 2x^2 -2x -4 > 0 So : 2(x-2)(x+1)>0 What is the next step?

OpenStudy (osanseviero):

Either one or the other is bigger than 0, so: x>2 U x>-1 (???)

terenzreignz (terenzreignz):

No..., actually, x > -1 doesn't do it justice, since -0.5 > -1 but if you plug in -0.5, it doesn't work...

OpenStudy (j2lie):

I think you distribute 2(x-2) and then you distribute 2(x+1).

terenzreignz (terenzreignz):

Nah... you just have to make sure both factors have the same sign~ that ensures that the resulting product is positive.

OpenStudy (j2lie):

how do you solve this problem? I learned inequality but I never learned this.

OpenStudy (jdoe0001):

@osanseviero your steps look right as terenzreignz said though, the -1 may just be an extraneous answer

terenzreignz (terenzreignz):

Here, let's make a table: Involving the 'critical values' of the left-part. Critical values are those values which would make the left-part either equal to zero or undefined. Thus, the critical values are 2 and -1 as you said. \[\Large \left.\begin{matrix}&&-\infty&-1&&2&+\infty\\x-2&&-&-&-&0&+\\x+1&&-&0&+&+&+\\(x-2)(x+1)&&+&0&-&0&+\end{matrix}\right.\] Here's a crude representation...

OpenStudy (j2lie):

I have never seen that.

OpenStudy (osanseviero):

Ok...I knew this method, I wanted to do algebraic

OpenStudy (osanseviero):

But thanks

terenzreignz (terenzreignz):

Algebraic is a bloody process, and it's not nearly as useful when you have more than two factors, just saying ^_^

OpenStudy (j2lie):

so you factor it?

terenzreignz (terenzreignz):

Conveniently, it is already factored j2lie :3

OpenStudy (osanseviero):

Ok, the table is done. Thanks :)

OpenStudy (j2lie):

don't you have to distribute the 2 to each terms in the parentheses?

OpenStudy (osanseviero):

Nope, it is an extra factor. The answer will be the same

OpenStudy (osanseviero):

It was actually with the two, I took it out to divide in different components

OpenStudy (j2lie):

so the 2 is not needed?

terenzreignz (terenzreignz):

Well, we want the left-part to be positive, but the 2 has no effect on that, since multiplying 2 to anything will not change the sign. That's why I left the 2 out of the table. ^_^

terenzreignz (terenzreignz):

But j2lie, if you insist... \[\Large \left.\begin{matrix}&&-\infty&-1&&2&+\infty\\2&&+&+&+&+&+\\x-2&&-&-&-&0&+\\x+1&&-&0&+&+&+\\(x-2)(x+1)&&+&0&-&0&+\end{matrix}\right.\]

OpenStudy (j2lie):

I have no idea why you're talking about. I haven't learn this.

terenzreignz (terenzreignz):

Oops... Anyway, j2lie, just multiply the coloured signs to get the sign at the bottom row, which is what we're looking for:\[\Large \left.\begin{matrix}&&-\infty&-1&&2&+\infty\\2&&\color{blue}+&\color{blue}+&\color{blue}+&\color{blue}+&\color{blue}+\\x-2&&\color{red}-&\color{red}-&\color{red}-&\color{green}0&\color{blue}+\\x+1&&\color{red}-&\color{green}0&\color{blue}+&\color{blue}+&\color{blue}+\\2(x-2)(x+1)&&+&0&-&0&+\end{matrix}\right.\]

OpenStudy (j2lie):

what do I do with the numbers?

terenzreignz (terenzreignz):

What numbers? The ones on top? Those are just markers. For instance, it shows that 2 is positive everywhere (obviously) x-2 is negative from negative infinity up to 2, where it is zero. From 2 onwards, it's positive. x+1 is negative from negative infinity up to -1, where it is zero. From -1 onwards, it's positive. With these knowledge, we can multiply these three to get the sign of the entire expression 2(x-2)(x+1) And we want to know at exactly which interval it is positive, which is now shown on the table ^_^

OpenStudy (j2lie):

I still don't get it. I'm sorry I haven't learn it yet.

terenzreignz (terenzreignz):

For instance, if we wanted to know the interval where 2(x-2)(x+1) is negative instead, the table shows that it's the interval BETWEEN -1 and 2.

terenzreignz (terenzreignz):

Well don't worry about it, j2lie, just appreciate it ^_^

OpenStudy (j2lie):

The interval is positive?

terenzreignz (terenzreignz):

The expression is positive at the intervals from negative infinity to -1, and from 2 to positive infinity.

OpenStudy (j2lie):

Yay I got it right.

OpenStudy (j2lie):

Thank you @terenzreignz.